48 students participated in evaluating small package of skittles. The following table shows the result of this study. AssumeQuestion 6) Company claims that more than 19% skittles are green. Do you agree with this claim? Use α=0.05.

Will rate and comment. Thank you.
Answer #1

\hat{p} = 576 / 2923 = 0.197

p = 0.19

1 - p = 0.81

The Hypothesis:

H0: p \leq 0.19

Ha: p > 0.19

This is a Right tailed Test.

The Test Statistic:

Z = \frac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}} = \frac{0.197-0.19}{\sqrt{\frac{0.19*0.81}{2923}}} = 0.96

The p Value:    The p value (Right tail) for Z = 0.96, is; p value = 0.1685

The Critical Value:   The critical value (Right tail) at \alpha = 0.05, Zcritical = +1.96

The Decision Rule: If Zobserved is > Zcritical Then Reject H0.

Also If the P value is < \alpha , Then Reject H0

The Decision:    Since Z observed (0.96) is < Zcritical (1.96), We Fail to Reject H0.

Also since P value (0.1685) is > \alpha (0.05), We Fail to Reject H0.

The Conclusion: There isn't sufficient evidence at the 95% significance level to conclude that the the number of green skittles is greater than 19%.

Therefore cannot agree with the company claim.